OpenStudy (anonymous):

The y-intercept for Q(x) = (x4 – 1)/(x2 – 4) is at (0, 1/4) True False

hartnn (hartnn):

to get y-intercept, you just put x=0 in your Q(x) so, plug in x=0 and tell me what u get ?

OpenStudy (anonymous):

lost :(

OpenStudy (anonymous):

help

hartnn (hartnn):

put x=0 in (x^4-1)/(x^2-4) so, whats (0^4-1)/(0^2-4) = ... ?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

?

hartnn (hartnn):

how did u get false ?

OpenStudy (anonymous):

because when i did the problem it was different

hartnn (hartnn):

see,0^4 = 0, 0^2 =0 so, you would get -1/-4 = 1/4 so your y -int is actually 1/4 so its true :)

OpenStudy (anonymous):

oooooohhhhh guess i didnt know how to do it thank you so much

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

do you think you could help me with this one am stuck

OpenStudy (anonymous):

What is the vertex of the following inequality: 2(2x2 – 3x) > -9

OpenStudy (anonymous):

What is the vertex of the following inequality: 2(2x^2 – 3x) > -9

OpenStudy (anonymous):

i have to pick one of them for the answer (-3/4, 27/4) (3/4, -27/4) (-3/4, -27/4) (3/4, 27/4)

hartnn (hartnn):

can i request you to ask only one question in one post ? close this and make a new post

OpenStudy (anonymous):

ok